DivineSolar (divinesolar):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

DivineSolar (divinesolar):

Anyone?

OpenStudy (mrnood):

OK a line with slope 4 will always be of the form y= 4x +b you know that on THIS line y= -1, when x = -3 so in the equation -1 = 4(-3) +b so find the value for b and oyu have your answer..

DivineSolar (divinesolar):

How do i solve for b?

OpenStudy (mrnood):

-1 = 4(-3) +b first expand the bracket 4 (-3)=?

DivineSolar (divinesolar):

-12?

OpenStudy (mrnood):

ok good so now the equation is -1 = -12 + b what is b?

DivineSolar (divinesolar):

@MrNood Im not sure to be honest.

OpenStudy (mrnood):

Well I can see that you are well beyond doing simple arithmetic - the 'straight line' stuff requires oyu to be able to add and subtract. Add 12 to each side of the equation. So for you have not really done anything towards this question. I am sure that you have enough maths behind you to solve this if you try Add 12 to each side of the equation: -1 = -12 + b

DivineSolar (divinesolar):

-13 = 1 + b?

DivineSolar (divinesolar):

Oh no wait its 11

DivineSolar (divinesolar):

So the answer is either A or B

DivineSolar (divinesolar):

Im thinking B

OpenStudy (mrnood):

OK - so you have the left side correct = 11 have a nother look at the right side (add 12)

DivineSolar (divinesolar):

-12 would become 0 wouldnt it.

OpenStudy (mrnood):

yes

DivineSolar (divinesolar):

so 11 = b

OpenStudy (mrnood):

yes and y=4x+b

DivineSolar (divinesolar):

Answer confirmed: y = 4x + 11

DivineSolar (divinesolar):

Correct?

OpenStudy (mrnood):

you know whether it is correct or not - you must have the confidence to submit. In the real world there is not always someone to confirm or 'mark' your work. Bridges may stand or fall by your calculations, or millions$ may be made or lost

DivineSolar (divinesolar):

Nice quote.